OpenStudy (anonymous):

On the first 4 quizzes of her math class, Ashley got an average score of 79. What does she need on the next quiz to have an overall average of 81?

OpenStudy (anonymous):

SOMEONE SPOTTED MY QUESTION!

OpenStudy (anonymous):

I posted it not too long ago, but it's still way far down there.

OpenStudy (kropot72):

The total score on the first 4 quizzes is given by multiplying the average score by 4: Total score for first 4 quizzes = 79 * 4 Total score needed for first 5 quizzes = 81 * 5 Score needed for the 5th quiz = (81 * 5) - (79 * 4) = you can calculate.

OpenStudy (anonymous):

Thanks. can you help me with a few more?